OpenStudy (anonymous):

Fill in the missing number 0.18 = ?/50

OpenStudy (anonymous):

9

OpenStudy (anonymous):

how

OpenStudy (anonymous):

.18=x/50 x=.18*50=9

OpenStudy (anonymous):

Write 0.25 as a fraction in simplest form. ?

OpenStudy (anonymous):

0.21/100=21/100

OpenStudy (anonymous):

awww

OpenStudy (anonymous):

25/100=1/4 :)

OpenStudy (anonymous):

so would 0.72 be 72/100?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you round it or anything

OpenStudy (anonymous):

reduce it further to simplest form

OpenStudy (anonymous):

how

OpenStudy (anonymous):

72/100(2 common) 36/50(2 common) 18/25

OpenStudy (anonymous):

Write 4.64 as an improper fraction in simplest form.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

??